OpenStudy (ujjwal):

The sum of infinite series \(1-3x+5x^2-7x^3+....to \infty\) when |x|<1 is: a)\[\frac{1-x}{1+x^2}\] b)\[\frac{1+x}{1-x^2}\] c)\[\frac{1-x}{(1+x)^2}\] d)\[\frac{1-x}{1-x^2}\]

OpenStudy (experimentx):

the pattern looks like -3, 5, -7, 9, - 11 ... if you expand the denominator using binomial theorem, ,,, you might get it.

OpenStudy (anonymous):

wait

OpenStudy (ujjwal):

I got this weird idea.. I arbitrarily put the value of x to be 0.1 and calculated up to 9x^4 (after that the number would get much smaller) which gave me a value 0.7439 and then i substituted x=0.1 in all options and BINGO!! i got 0.7438 in option C.. So, C seems to be correct!!

OpenStudy (experimentx):

well ... that would work out too. i was guessing 1/(1+x^2) = 1 + ..x^2 + ..x^4 + ..x^6 + .. (1+x)(1 + ..x^2 + ..x^4 + ..x^6 + ..) = 1 + x + .a.x^2 + .a. x^3 + ... so it couldn't be 1/(x^2+ 1)

OpenStudy (anonymous):

The answer is CCCCCCCCCCCCC

OpenStudy (experimentx):

Woops \[ \frac{1-x}{1-x^2} = {1 \over 1+x} = {1 \over 1 -(-x)} = 1 - x + x^2 + x^3 - ... \]

OpenStudy (experimentx):

similarly \[ \frac{1+x}{1-x^2} = 1 +x +x^2 + x^3 + ... \]

OpenStudy (experimentx):

for this \[ \frac{1-x}{1+x^2} \] you should have alternating pattern ... which is not in the the series. so generating function for series is C

OpenStudy (experimentx):

Woops!! error \[ \frac{1-x}{1-x^2} = {1 \over 1+x} = {1 \over 1 -(-x)} = 1 - x + x^2 - x^3 + ...\]

hartnn (hartnn):

this is arithmetico geometric series...u just need a formula.

OpenStudy (anonymous):

Ok here it .... Let s =1-x+3x^2-5x^3+.....(+-)(2n-1)x^n Now let me ask what is x^2*s

OpenStudy (experimentx):

seems that way too ... though deriving the left from right is quite difficult.

OpenStudy (anonymous):

And what did you get for the expression s-x^2s

hartnn (hartnn):

or @Zekarias method is also good.

OpenStudy (anonymous):

thus s(1-x^2)=1-3x+4(x^2-x^3+x^4-.........)

OpenStudy (anonymous):

Will you finish then?

OpenStudy (ujjwal):

Thanks everybody!!